OpenStudy (anonymous):

Nicole bought 8 CDs. a week later. ome fourth of them are destroyed. there are now only 3 left. With how many did she sart with? What's the equation?

hero (hero):

(x+8)/4 = 3

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

@Hero One fourth of them are destroyed

hero (hero):

that's what the division by 4 is for

OpenStudy (anonymous):

your name reminds me of Laimoonah lol

OpenStudy (anonymous):

Hahah that rhymes of course.

OpenStudy (anonymous):

(x+8)/4 = number of CDs destroyed

OpenStudy (anonymous):

Haha its a nice name i love it

OpenStudy (anonymous):

So who's correct @monybird or @Hero

OpenStudy (anonymous):

Thanks!

hero (hero):

so she had 12 to begin with

hero (hero):

because 12/4 = 3

OpenStudy (anonymous):

Ok.. Thank you

hero (hero):

Sorry moneybird, you're wrong on this one. But you're really a really smart person without question

OpenStudy (anonymous):

Ya i think im wrong sorry

OpenStudy (anonymous):

Its Okay